A woman makes $6 an hour and $9 an hour in overtime. She made $324 clocking 50 hours. How many hours were overtime.
Answer by richwmiller(17219)

r+o=50,
6*r+9*o=324
r=50-o
We have two equations with two variables.
We will substitute to make one equation with one variable
6*(50-o)+9*o=324
300-6o+9*o=324
3*o=24
o=8
r=50-o
r=42 at $6 and o=8 at $9
check
42+8=50
6*r+9*o=324
6*42+9*8=324
252+72=324
324=324
ok
codetickv